A 1:1 mixture by moles of nitrous oxide and oxygen is often used as a sedative in dentistry. If the total pressure of this mixture in a cylinder is 2.50 atm, what is the partial pressure of each gas?
Answer #1

mol of N2O = 1 mol

mol of O2 = 1 mol

Total number of mol = n1+n2

= 1 + 1

= 2 mol

Partial pressure of each components are

p(N2O),p1 = (n1*Ptotal)/total mol

= (1 * 2.5)/2

= 1.25 atm

p(O2),p2 = (n2*Ptotal)/total mol

= (1 * 2.5)/2

= 1.25 atm

Answer:

partial pressure of each gas is 1.25 atm